# Negative Yields Positive: Unified Dual-Path Adapter for Vision-Language Models

### v1

@@ Abstract
Recently, large-scale pre-trained Vision-Language Models (VLMs) have demonstrated great potential in learning open-world visual representations, and exhibit remarkable performance across a wide range of downstream tasks through efficient fine-tuning. In this work, we innovatively introduce the concept of dual learning into fine-tuning VLMs, i.e., we not only learn what an image is, but also what an image isn’t. Building on this concept, we introduce a novel DualAdapter approach to enable dual-path adaptation of VLMs from both positive and negative perspectives with only limited annotated samples. In the inference stage, our DualAdapter performs unified predictions by simultaneously conducting complementary positive selection and negative exclusion across target classes, thereby enhancing the overall recognition accuracy of VLMs in downstream tasks. Our extensive experimental results across 15 datasets validate that the proposed DualAdapter outperforms existing state-of-the-art methods on both few-shot learning and domain generalization tasks while achieving competitive computational efficiency. Code is available at https://github.com/zhangce01/DualAdapter.
Keywords:
Vision-Language Models CLIP Few-Shot Classification Domain Generalization Transfer Learning
@@ 1 Introduction
Large-scale pre-trained Vision-Language Models (VLMs), such as CLIP [48] and ALIGN [24], provide a new paradigm for multi-modal learning and open-world visual recognition [78, 8]. Recently, researchers have demonstrated the robustness and effectiveness of these VLMs in interpreting complex visual and textual inputs, as evidenced by their superior performance in a variety of vision-language tasks, e.g., image captioning [73, 21, 60], visual question answering [52, 25, 31], and other domains [7, 85, 54, 58].
Current VLMs excel in zero-shot inference [86, 15], but struggle in transferring to specific downstream tasks, primarily due to domain shifts and mismatches between pre-training and task-specific data [10, 42]. Additionally, the extensive number of parameters and demanding computational costs make it intractable for individual practitioners to re-train or fine-tune the entire models to enhance their task-specific performance [84, 12]. To address this, recent studies have focused on designing efficient fine-tuning strategies that maintain the core pre-trained encoders frozen while incorporating lightweight adapter [12, 79] modules or optimizing prompts[84, 83]. These methods enable effective task-specific knowledge extraction from a limited amount of annotated training samples, on top of the inherited foundational knowledge from pre-training.
Figure 1: Dual-path inference examples. We show four sample images from diverse recognition tasks, along with their similarities (scaled by 100) to positive/negative text descriptions after few-shot dual-path adaptation. Intuitively, an image is more likely to be classified into a specific class if it exhibits higher positive similarity and lower negative similarity. We can observe that while the positive branch alone may struggle to distinguish among some closely related classes, incorporating the negative path, which eliminates certain incorrect classes, enhances its ability to accurately identify the true class.
In this work, we introduce the innovative concept of dual inference in the context of fine-tuning VLMs. Take, for instance, a basic binary pet classification task: distinguishing cats from dogs. To identify an image as a cat, we can approach the task from both the positive and negative perspectives: one might directly identify the image as a cat (selection), or alternatively, one can also first determine that it is not a dog (exclusion), thereby indirectly concluding it’s a cat. These approaches, namely, positive selection and negative exclusion, complement each other and collectively enhance the overall classification accuracy in general classification tasks.
Building on this concept, we introduce a novel DualAdapter approach for few-shot adaptation of VLMs that conducts positive selection and negative exclusion across a set of target classes. While CLIP typically performs classification tasks using prompts such as “A photo of a {CLASS}”, we recognize its ability to understand negation, employing prompts like “A photo of no {CLASS}”. Utilizing this capability, we design our DualAdapter to adapt VLMs from a dual perspective: we propose to learn not only what an image is but also what an image isn’t by matching the input image feature with the positive and negative text features. Intuitively, our approach first determines a set of likely candidate classes through image classification but, in a second step, refines these classification logits by asking the inverse question of what an image isn’t across the entire set of target classes. In Fig. 1, we showcase several dual-path inference examples to illustrate that when positive selection struggles with distinguishing between similar classes, employing negative exclusion can provide further insights to improve recognition accuracy.
Moreover, in DualAdapter, we symmetrically apply dual-path adaptation to the visual modality, treating image features of a specific class as positive and those from all other classes as negative. Recognizing the variability in image quality and representativeness, where some image features are noisy and not equally representative of their respective classes, we implement an unsupervised similarity-based label refinement technique. This method assigns confidence scores to each image, reducing the impact of outliers or less representative samples during few-shot adaptation, ensuring more accurate and reliable classification.
@@
To empirically validate DualAdapter, we perform thorough evaluations on few-shot learning and domain generalization tasks using 15 diverse recognition datasets. These experimental results demonstrate the effectiveness of DualAdapter in adapting VLMs for downstream tasks and verify its superior robustness to distribution shifts. Additionally, our efficiency benchmarks indicate that DualAdapter achieves competitive efficiency compared to other state-of-the-art methods.
In summary, our primary contributions are listed as follows:
  • $\bullet$
    We explore and exploit the negative inference capabilities of VLMs, and introduce a unique dual-path adaptation approach (i.e., DualAdapter) for CLIP.
  • $\bullet$
    To mitigate the impact of noisy samples in few-shot adaptation, we introduce a novel unsupervised similarity-based label refinement technique.
  • $\bullet$
@@
    Through extensive experiments, we demonstrate that our DualAdapter outperforms other state-of-the-art methods in both adaptation performance and generalizability across 15 diverse recognition datasets.
@@ 2 Related Work
Vision-Language Models. In recent years, extensive Vision-Language Models (VLMs) have been developed to bridge the vision and language modalities through large-scale pre-training [48, 24]. Generally, these pre-trained VLMs can be classified into four categories based on their pre-training objectives: masked language modeling [37, 30, 27], masked region prediction [37, 30, 59], image-text matching [30, 27, 61, 31], and contrastive learning [48, 24, 74, 77, 33]. Researchers have demonstrated that these pre-trained VLMs exhibit remarkable open-world visual understanding capabilities [8], achieving state-of-the-art performance across a range of open-world tasks, such as zero-shot learning [48, 51, 41], open-world segmentation [34, 72, 49], and open-world detection [53, 16, 9]. In this work, we mainly focus on contrastive-based VLMs, especially CLIP [48], which has become the current mainstream approach in this field. Nonetheless, we believe that the core concept of this work is adaptable and can be extended to other VLMs as well.
@@
Efficient Fine-Tuning for VLMs. Given the substantial size of the VLMs, recent research efforts [84, 12, 79, 80] are focusing on the development of lightweight fine-tuning techniques to efficiently adapt VLMs for downstream visual tasks. These methods can generally be divided into two primary categories: prompt-based learning and adapter-style fine-tuning. Specifically, prompt-based learning methods [84, 83, 2, 62] aim to optimize the input prompts from downstream data, while adapter-style fine-tuning approaches [12, 79, 64, 75, 88] directly tune the extracted visual and textual representations. Our proposed DualAdapter follows the line of adapter-style fine-tuning, but it distinguishes itself from previous work [79, 75] by incorporating negative adapters for the first time. Moreover, we empirically demonstrate that leveraging negative cues from CLIP can effectively improve both few-shot classification performance and generalization capability.
@@
Negative Learning. Given that obtaining typical positive labels (which indicate the categories an image belongs to) can be costly and labor-intensive, researchers have proposed an alternative approach known as the indirect negative learning [23, 76] paradigm. This method focuses on learning from easily accessible negative/complementary labels, which specify the categories to which an image does not belong. In recent years, negative learning has been effectively applied to various vision applications, such as image recognition [23, 28, 13], few-shot learning [22, 69], and semantic segmentation [67, 47]. In this work, empowered by the strong image-text association capabilities of the advanced pre-trained CLIP [48] model, we enable negative learning within CLIP without the need for explicit negative labels. Instead, we design negative prompts and negative image samples to guide the CLIP model to predict from a negative perspective, which in turn enhances its overall positive prediction capability. To the best of our knowledge, there are only two concurrent approaches (i.e., CLIPN [66] and LSN [43]) that work in a similar manner: they both enable the negative predictions of CLIP to identify unknown out-of-distribution samples. However, we formulate negative learning from a different and more fundamental perspective, i.e., using negative prediction to directly improve the accuracy of positive prediction.
Contrastive Learning. We also recognize the underlying concept behind our proposed DualAdapter has parallels with existing supervised contrastive learning literature [14, 82, 26]: Our method also aims to maximize similarities between positive image-text/image-image pairs, while simultaneously minimizing similarities between negative pairs during few-shot adaptation. However, our work focuses on a fundamentally different dual learning perspective, i.e., we explicitly train two distinct positive and negative classifiers for ensembling using a single image input at a time, rather than applying supervised contrastive loss to train a single classifier with image pairs as inputs. It is also important to note that contrastive learning is not directly applicable in our fine-tuning setting because we keep the CLIP image encoder frozen. Thereby, the extracted image features are fixed and cannot be pulled together or pushed apart.
@@
Figure 2: An overview of our proposed DualAdapter. We introduce four adapters, each of which builds a cache with positive/negative text/image features that are updated during few-shot training. Given an image to be classified, the classification logit for a specific class increases when the image feature closely aligns with the corresponding features in the positive cache and diverges from those in the negative cache.
3 DualAdapter: Unified Dual-Path Adaptation of VLMs
We introduce a novel framework DualAdapter, as illustrated in Fig. 2, to enable efficient dual-path adaptation of VLMs. Our DualAdapter consists of four adapters: two that positively adapt CLIP to enhance its ability to accurately identify the true class of an input image, and two that negatively adapt CLIP to effectively exclude incorrect candidate classes. Furthermore, we formulate few-shot learning as a noisy problem, and propose an unsupervised similarity-based label refinement approach to select high-quality representative samples to more effectively adapt CLIP to downstream tasks. In the following sections, we first briefly review zero-shot CLIP, then introduce each component of our DualAdapter in detail.
@@ 3.1 Preliminary: A Revisit of CLIP
By leveraging contrastive training on billions of image-text pairs from the Internet, CLIP has shown an outstanding ability to align vision and language modalities [48]. In this work, we employ CLIP’s pre-trained visual encoder $\mathcal{F}_{\mathsf{V}}$ and textual encoder $\mathcal{F}_{\mathsf{T}}$ to map the images and textual descriptions into a shared $d$-dimensional embedding space. For a $C$-class classification task, CLIP performs zero-shot predictions by assessing the similarity between the image feature and the text feature specific to each class $c\in\{1,...,C\}$ as follows:
@@
$f_{\mathcal{I}}=\mathcal{F}_{\mathsf{V}}(\mathcal{I}),\quad f_{t_{c}}=\mathcal{F}_{\mathsf{T}}(\mathcal{T}_{c}),\quad\mathbb{P}(y=c|\mathcal{I})=\frac{\exp\left(\mathrm{cos}\left(f_{t_{c}},f_{\mathcal{I}}\right)/t\right)}{\sum\nolimits_{t^{\prime}}{\exp\left(\mathrm{cos}\left(f_{t^{\prime}},f_{\mathcal{I}}\right)/t\right)}},$ (1)
[…]
@@
To streamline this process, a weight matrix can be precomputed and stored in a textual cache, which concatenates the textual features associated with each class, denoted as $\mathsf{T}_{\mathtt{cache}}=[f_{t_{1}}\,f_{t_{2}}\,\cdots\,f_{t_{C}}]^{\top}\in\mathbb{R}^{C\times d}$. Subsequently, we can efficiently obtain the logit $\mathcal{S}$ and the final prediction $\mathbb{P}(y|\mathcal{I})$ via vectorized computation:
@@
$\mathcal{S}=f_{\mathcal{I}}\mathsf{T}_{\mathtt{cache}}^{\quad\,\,\,\,\top}\in\mathbb{R}^{1\times C},\quad\mathbb{P}(y|\mathcal{I})=\mathsf{Softmax}\left(\mathcal{S}\right).$ (2)
3.2 Positive Adaptation
Although the zero-shot CLIP model is trained on web-scale datasets and encodes extensive prior knowledge, it cannot be effectively applied to downstream tasks due to the presence of domain shifts and a deficiency in task-specific knowledge. To address this, we efficiently leverage few-shot annotated samples to distill new, task-specific knowledge. As we discussed in Section 1, our DualAdapter enables few-shot adaptation from both positive and negative perspectives. In this section, we first focus on the positive side, i.e., refining CLIP’s ability to more precisely predict the true class of the input image in a specific downstream task.

### v2

# Enhancing Vision-Language Few-Shot Adaptation with Negative Learning

@@ Abstract
Large-scale pre-trained Vision-Language Models (VLMs) have exhibited impressive zero-shot performance and transferability, allowing them to adapt to downstream tasks in a data-efficient manner. However, when only a few labeled samples are available, adapting VLMs to distinguish subtle differences between similar classes in specific downstream tasks remains challenging. In this work, we propose a Simple yet effective Negative Learning approach, SimNL, to more efficiently exploit the task-specific knowledge from few-shot labeled samples. Unlike previous methods that focus on identifying a set of representative positive features defining “what is a {CLASS}", SimNL discovers a complementary set of negative features that define “what is not a {CLASS}", providing additional insights that supplement the positive features to enhance task-specific recognition capability. Further, we identify that current adaptation approaches are particularly vulnerable to potential noise in the few-shot sample set. To mitigate this issue, we introduce a plug-and-play few-shot instance reweighting technique to suppress noisy outliers and amplify clean samples for more stable adaptation. Our extensive experimental results across 15 datasets validate that the proposed SimNL outperforms existing state-of-the-art methods on both few-shot learning and domain generalization tasks while achieving competitive computational efficiency. Code is available at https://github.com/zhangce01/SimNL.
@@ 1 Introduction
Over the past decade, deep learning models have achieved remarkable progress in various vision tasks [41, 42], largely due to training on extensive supervised datasets [6, 27]. However, in real-world scenarios, acquiring such large-scale datasets in specific domains (e.g., satellite and aircraft images) is often impractical due to the time-consuming and costly nature of the collection and annotation process. As a result, there is often not enough data in these domains to capture the variability of the novel classes involved, making it challenging to develop robust models from scratch that can effectively generalize to unseen data [58, 47].
Recent advances in Vision-Language Models (VLMs), such as CLIP [38] and ALIGN [21], provide a promising alternative approach for building robust models in a low-data regime. Specifically, by utilizing large-scale pre-training on web-scale datasets, these VLMs have demonstrated remarkable zero-shot performance and transferability [38, 64, 67, 26], which enables the adaptation of these models to downstream tasks in a data-efficient manner. Currently, researchers have developed two primary few-shot adaptation strategies for VLMs: prompt-based learning [74, 73, 45, 50, 72] and adapter-style fine-tuning [10, 71, 25, 65, 70], to enable effective task-specific knowledge extraction. However, due to the very limited number of annotated samples available, these methods still struggle to discern subtle differences between similar classes in specific downstream tasks.
Figure 1: Negative learning provides complementary information for more accurate recognition. (Left) Grad-CAM [44] visualization of the similarity heatmaps with the learned positive and negative features of the ground-truth class; (Middle) Similarities (scaled by 100) to the learned positive and negative features of five similar candidate classes. While the positive branch alone may fail to distinguish among some closely related classes, incorporating the negative classifier, which eliminates certain incorrect classes, enhances the model’s ability to accurately identify the true class; (Right) Performance comparisons with other state-of-the-art methods in 16-shot scenarios.
In this work, inspired by the negative learning literature [20, 66, 49], we propose a simple yet effective negative learning approach, SimNL, for more effectively adapting VLMs to downstream tasks. Specifically, while previous adaptation methods [74, 73, 71] typically focus on identifying a set of representative features that define “what is a {CLASS}" during few-shot adaptation, we propose to discover a complementary set of negative features that define “what is not a {CLASS}" to better exploit the limited task-specific knowledge from the few-shot samples. As shown in Figure 1 (Left), this complementary set of negative features guides our model to pay attention to more diverse attributes when classifying an image, evaluating both the presence of class-specific characteristics and the absence of characteristics not associated with the class. Building on these discovered features, our SimNL framework incorporates two coordinated classifiers: one performing similarity matching with the positive features and the other performing dissimilarity matching with the negative features. In Figure 1 (Middle), we illustrate that when the positive classifier struggles with distinguishing between similar classes, employing the negative one can provide further insights to improve recognition accuracy. The performance comparisons in Figure 1 (Right) further highlight that our SimNL method, by incorporating an additional negative classifier, consistently achieves superior few-shot adaptation performance across 11 various datasets when compared to existing state-of-the-art approaches.
Another important challenge often overlooked in existing adaptation methods is the presence of noise in the few-shot sample set. Most existing approaches assume that each few-shot sample is carefully curated to accurately represent its class, but such assumptions rarely holds in practice. Given the extremely small size of the few-shot sample set, the models trained from these samples are particularly vulnerable to noise. To address this important issue, we extend SimNL by reweighting each few-shot instance with a confidence score, which suppresses outliers and amplifies clean samples to ensure stable adaptation. This plug-and-play technique can also be applied to other adapter-style fine-tuning methods, such as Tip-Adapter-F [71], making them more robust to noisy data.
@@
To empirically validate the effectiveness of our SimNL, we conduct comprehensive evaluations on few-shot learning and domain generalization tasks across 15 diverse recognition datasets. These experimental results demonstrate the effectiveness of SimNL in adapting VLMs for downstream tasks and verify its superior robustness to distribution shifts. We also demonstrate that our proposed instance reweighting significantly enhances the model’s robustness to label noises in the few-shot sample set. Additionally, SimNL also exhibits competitive efficiency.
Our primary contributions are summarized as follows:
  • We propose a simple yet effective negative learning approach, i.e., SimNL, to efficiently adapt CLIP to downstream tasks. Specifically, SimNL introduces an innovative application of negative learning to adapter-style fine-tuning for few-shot adaptation of VLMs.
  • To mitigate the impact of noisy samples in few-shot adaptation, we introduce a plug-and-play few-shot instance reweighting technique that assigns non-uniform confidence scores to each sample.
@@
  • Through extensive experiments, we demonstrate that our SimNL consistently outperforms other state-of-the-art methods across 15 diverse recognition datasets.
@@ 2 Related Work
Efficient Adaptation for VLMs. In recent years, extensive Vision-Language Models (VLMs) have been developed to bridge the vision and language modalities through large-scale pre-training [38, 21]. Given the substantial size of the VLMs, recent research efforts [74, 10, 71, 69] are focusing on the development of lightweight fine-tuning techniques to efficiently adapt VLMs for downstream visual tasks. These methods can generally be divided into two categories: prompt-based learning and adapter-style fine-tuning. Specifically, prompt-based learning methods [74, 73, 4, 50, 61] aim to optimize the input prompts from downstream data, while adapter-style fine-tuning approaches [10, 71, 52, 65, 76] directly tune the extracted visual and textual representations. In this work, we aim to enhance adapter-style fine-tuning by incorporating negative learning into vision-language few-shot adaptation. We also empirically validate that leveraging negative cues from CLIP can effectively improve both few-shot classification performance and generalization capability.
Few-Shot Learning. Few-shot learning aims to quickly adapt a model to new categories using only a few examples. Traditional few-shot learning methods primarily rely on meta-learning and can be roughly separated into two groups: metric-based methods [54, 46, 68] and optimization-based methods [39, 43, 12]. However, these methods depend on training with base datasets, limiting their applicability in real-world scenarios. Recent developments in large-scale pre-trained VLMs [38, 1] present a promising alternative due to their exceptional zero-shot capabilities. Researchers have demonstrated that with efficient adaptation, VLMs can also excel in few-shot learning tasks [74, 73, 71, 17].
@@
Negative Learning. Given that obtaining typical positive labels (which indicate the categories an image belongs to) can be costly and labor-intensive, researchers have proposed an alternative approach known as the indirect negative learning [20, 66] paradigm. This method focuses on learning from negative/complementary labels, which specify the categories to which an image does not belong. In recent years, negative learning has been effectively applied to various vision applications, such as image recognition [20, 23, 11], few-shot learning [18, 60], and semantic segmentation [57, 37]. In this work, empowered by the strong negative understanding capabilities of the large-scale pre-trained CLIP model [56, 32, 19], we enable an alternative approach to negative learning, where we explicitly train a negative classifier without the need for complementary labels. We provide more discussions on the differences between our method and traditional negative learning in Section B.1. We also note that recent work on TDA [22] shares a similar approach to ours, as both methods incorporate negative caches to enhance the generalizability of VLMs. However, their method focuses on test-time adaptation, where training data is unavailable, making it unsuitable for direct application in few-shot adaptation scenarios.
3 Method
In this section, we introduce SimNL, a novel approach to vision-language few-shot adaptation, as shown in Figure 2.
@@ 3 Method
Figure 2: An overview of our proposed SimNL. We construct and learn the positive and negative CLIP-based classifiers across visual and textual modalities. Given an image to be classified, the classification logit for a specific class increases when the image feature $f_{v}$ closely aligns with the corresponding positive features $f_{t}^{+},f_{v}^{+}$ and diverges from negative features $f_{t}^{-},f_{v}^{-}$.
3.1 Problem Formulation
We consider the problem of $C$-way $K$-shot few-shot classification problem, in which only $K$ labeled examples are provided for each of the $C$ classes. Let the feature space be $\mathcal{X}$ and the label space be $\mathcal{Y}=\{1,...,C\}$, the few-shot instance $\boldsymbol{x}\in\mathcal{X}$ and its labels $y\in\mathcal{Y}$ are sampled from probability distribution $P(\boldsymbol{x},y)$. The corresponding one-hot encoded labels are represented by $\boldsymbol{y}\in\{0,1\}^{C}$.
Positive Learning. In this study, we employ a CLIP-based classifier with parameters $\theta$, denoted as $\mathcal{F}_{\theta}:\mathcal{X}\rightarrow\mathbb{R}^{C}$, to project each input into a $C$-dimensional score space. The classification probabilities are obtained using $\boldsymbol{p}=\texttt{Softmax}\left(\mathcal{F}_{\theta}(\boldsymbol{x})\right)\in\Delta^{C-1}$, where $\Delta^{C-1}$ represents the $(C-1)$-dimensional probability simplex. Our goal is to learn a robust classifier $\mathcal{F}_{\theta}$ that minimizes the expected risk; that is,
$\min_{\theta}\mathcal{R}(\mathcal{F}_{\theta})=\mathbb{E}_{(\boldsymbol{x},y)\sim P(\boldsymbol{x},y)}\left[\mathcal{L}\left(\mathcal{F}_{\theta}(\boldsymbol{x}),y\right)\right],$
$\text{where}\,\,\mathcal{L}\left(\mathcal{F}_{\theta}(\boldsymbol{x}),y\right)=-\sum_{k=1}^{C}\boldsymbol{y}_{k}\log\boldsymbol{p}_{k},$ (1)
where $\boldsymbol{p}_{k}$ denotes the probability for the $k$-th class and $\mathcal{L}\left(\mathcal{F}_{\theta}(\boldsymbol{x}),y\right)$ computes the cross-entropy loss. We aim to maximize the probability $\boldsymbol{p}_{y}\rightarrow 1$ for the true label.
Negative Learning. In this work, we apply the concept of negative learning to vision-language few-shot adaptation by employing a distinct CLIP-based negative classifier $\mathcal{G}_{\varphi}:\mathcal{X}\rightarrow\mathbb{R}^{C}$ with parameters $\varphi$. This classifier predicts the negative probability $\bar{\boldsymbol{p}}=\texttt{Softmax}(\mathcal{G}_{\varphi}(\boldsymbol{x}))$ that the image does not belong to specific classes. The expected classification risk for $\mathcal{G}_{\varphi}$ can be written as
$\min_{\varphi}\mathcal{R}(\mathcal{G}_{\varphi})=\mathbb{E}_{(\boldsymbol{x},y)\sim P(\boldsymbol{x},y)}\left[\mathcal{L}\left(\mathcal{G}_{\varphi}(\boldsymbol{x}),y\right)\right],$
$\text{where}\,\,\mathcal{L}\left(\mathcal{G}_{\varphi}(\boldsymbol{x}),y\right)=-\sum_{k=1}^{C}\boldsymbol{y}_{k}\log(1-\bar{\boldsymbol{p}}_{k}).$ (2)
By optimizing this risk, we aim to reduce the negative probability $\bar{\boldsymbol{p}}_{y}\rightarrow 0$ for the true label.
Finally, we ensemble the optimized classifiers from both positive and negative learning to create an enhanced classifier $\mathcal{F}_{\theta}\oplus\mathcal{G}_{\varphi}:\mathcal{X}\rightarrow\mathbb{R}^{C}$ for testing. This is equivalent to combining the positive and negative predicted probabilities as $\hat{\boldsymbol{p}}=\lambda\boldsymbol{p}+(1-\lambda)(1-\bar{\boldsymbol{p}})$, where $\lambda$ is a balancing hyper-parameter.
Differences with Contrastive Learning. Note that the concept of “negative" differs in our negative learning approach compared to contrastive learning: (1) In our negative learning, “negative" specifically refers to a negative classifier. We explicitly train another negative classifier $\mathcal{G}_{\varphi}$ to ensemble with the positive classifier; (2) In contrastive learning, “negative" refers to the negative sample pairs that are constructed and utilized during training. In Appendix B.2, we show that the training objectives for both the positive and negative CLIP-based classifiers are contrastive in nature.
@@ 3.2 Preliminary: A Revisit of CLIP
In this work, our classifiers are based on CLIP’s pre-trained visual encoder $\mathcal{F}_{\mathsf{V}}$ and textual encoder $\mathcal{F}_{\mathsf{T}}$. Specifically, CLIP [38] performs zero-shot predictions by assessing the similarity between the image feature and the text feature specific to each class $c\in\{1,...,C\}$ as follows:
$f_{v}=\mathcal{F}_{\mathsf{V}}(\boldsymbol{x}),\quad f_{t_{c}}=\mathcal{F}_{\mathsf{T}}(\mathcal{T}_{c}),$
@@
$\mathbb{P}(y=c|\boldsymbol{x})=\frac{\exp\left(\mathrm{cos}\left(f_{t_{c}},f_{v}\right)/t\right)}{\sum\nolimits_{t^{\prime}}{\exp\left(\mathrm{cos}\left(f_{t^{\prime}},f_{v}\right)/t\right)}},$ (3)
@@
where $\boldsymbol{x}$ is the input image, and $\mathcal{T}_{c}$ represents the text description for class $c$ (e.g., “A photo of a {CLASSc}”). The parameter $t$ is the temperature hyperparameter, and $\mathrm{cos}\left(f_{t},f_{v}\right)=f_{v}^{\top}f_{t}$ computes the cosine similarity. To streamline this process, we can precompute a textual cache, which concatenates the textual features associated with each class, denoted as $\mathsf{T}_{\mathtt{cache}}=[f_{t_{1}}\,f_{t_{2}}\,\cdots\,f_{t_{C}}]^{\top}\in\mathbb{R}^{C\times d}$. Subsequently, we can obtain the final prediction $\mathbb{P}(y|\boldsymbol{x})$ via vectorized similarity matching:
@@
$\mathcal{S}=f_{v}\mathsf{T}_{\mathtt{cache}}^{\quad\,\,\,\,\top}\in\mathbb{R}^{C},\quad\mathbb{P}(y|\boldsymbol{x})=\mathtt{Softmax}\left(\mathcal{S}\right).$ (4)
3.3 Learning the Positive Classifier
In this section, we focus on the positive perspective, namely, constructing and optimizing a learnable CLIP-based positive classifier. This similarity-based classifier is designed to identify a set of representative features that enhance the accuracy of predicting the true class of the input image.